Under what condition will the magnitude of the torque acting on an electric dipole in a uniform electric field be equal to its potential energy


1
Expert's answer
2021-06-02T13:03:34-0400


Torque="\tau"

Electric dipole moment =p

Uniform electric field =E

Dipole moment p=q.d

Torque"\tau=p\times E"


"\tau =pEsin\theta"

"\theta=90°"

sin90°=1

"\tau=pE"

Energy

"U=-p.E"

"U=-pEcos\theta"

"\theta=180°"

Cos180°=-1

U=pE

Then

"\theta=90°" torque is equal to"\theta =180 °" electric potential energy